OpenStudy (anonymous):

What is the solution to the equation? 1/sqrt8 = 4^(m+3)? My answer - m=5/4

hartnn (hartnn):

-3/2 = 2(m+3) -3 = 4m+12 hmm.... how you got m=5/4 ??

OpenStudy (anonymous):

wow you have an "A" next to your name!!!!! okay I plugged it in and that's what I got. I knew it have to be positive.

hartnn (hartnn):

can you try solving it? 8 = 2^3 sqrt 8 = 2^{3/2} 4 = 2^2 plug these in!

OpenStudy (anonymous):

wait I did it a different way this time

OpenStudy (anonymous):

I got -15/4?

hartnn (hartnn):

thas correct! good :)

OpenStudy (anonymous):

really? yesss!!!!!!!!!! :)))))) thank you!!!!

hartnn (hartnn):

welcome ^_^ and thats A is because i am the Ambassador of Openstudy now :D

OpenStudy (anonymous):

oh my gosh congrats!!!!! you deserve it!!!!!! what does that mean for you now?

hartnn (hartnn):

thank you! :) that means i welcome new users here and make it easy to browse this site and make them feel good about coming here for getting and giving help :)

OpenStudy (anonymous):

That's so great of you! I wouldn't want it to be anyone else but you! Good luck ;)

hartnn (hartnn):

thanks again :)